OpenStudy (anonymous):

The marked price of a bag is $900. If the bag is sold at the marked price, then the percentage profit is 50%. If the bag is sold at a 20% discount on the marked price, then the profit is $___?

OpenStudy (mandre):

Oops. Let's try again.

OpenStudy (anonymous):

so messy...i didn't understand it..

OpenStudy (mandre):

x + 50% = $900

OpenStudy (anonymous):

OHH

OpenStudy (anonymous):

x+50%??

OpenStudy (mandre):

x * 1.5 = 900. Calculate x

OpenStudy (mandre):

To calculate the profit we need to find out how much the shopowner paid for it

OpenStudy (mandre):

That would be x

OpenStudy (anonymous):

x=300

OpenStudy (mandre):

300 + 50% = 450

OpenStudy (anonymous):

ohh, calculation mistake 1800

OpenStudy (mandre):

LOL. That would not be a profit would it?

OpenStudy (anonymous):

what?

OpenStudy (mandre):

Take some number and add half to it to get $900 so technically it is x + x + x = 900

OpenStudy (anonymous):

are you sure "+" ?

OpenStudy (mandre):

Yes 3x = 900

OpenStudy (anonymous):

x=300? -_-?

OpenStudy (mandre):

REmember we want 2 x

OpenStudy (anonymous):

600

OpenStudy (mandre):

600 + 50% = ?

OpenStudy (mandre):

Price = $900 - 20% New profit = Price - $600.

OpenStudy (anonymous):

Half of 900 is 450 thats how much it cost to make ice. 20 percent of 900 is 180 so at 20 percent discount ice sells for 620. 620 minus 450 is your profit

OpenStudy (anonymous):

600+50% ?

OpenStudy (anonymous):

Dios mio guys essy math

OpenStudy (mandre):

Really easy math. Except that 900-180 = 720 And if the cost price is $450 dollars then $900 would be 100% profit, not 50%.

OpenStudy (mandre):

kryton, what is 50% of 600?

OpenStudy (anonymous):

Selling price minus principal equals profit

OpenStudy (anonymous):

If your selling price is 900 and your profit is 50% of 900 then the principal is 450

OpenStudy (mandre):

My mistake then. You still got the other one wrong. $900-$180 = $720

OpenStudy (anonymous):

I did

OpenStudy (mandre):

We lost kryton along the way.

OpenStudy (anonymous):

720 minus 450 is the new profit at 20 percent off

OpenStudy (anonymous):

i have a doubt in this one. c.p+profit=s.p. now...its given that profit is 50% when sold at 900$.profit is always relative to c.p.so this means 50% profit=c.p/2.from this c.p+c.p/2=900. 3 x c.p/2=900.we get c.p=600 and profit as 300. if sold at 20% discount,s.p=720.then profit =720-600=120$ Am i doing it wrong?? @cheme @Mandre @kryton1212

OpenStudy (mandre):

cheme disagrees with you. I don't know financial nomenclature so I was winging it with plain math.

OpenStudy (mandre):

And I presumed if you paid 450 for something and sold it for 900 your profit would be 100%

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

@Luigi0210 @evilmath some help guys

OpenStudy (anonymous):

@RANE @Ryaan @Imtiaz7

OpenStudy (rane):

why is your q' closed?

OpenStudy (anonymous):

it cannot open again, never mind, just answer is ok

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

selling price=cost price +profit.. given that s.p=900$ and corresponding profit is 50%. profit means the additional money we make over a bargain

OpenStudy (anonymous):

@kryton1212 okay so far?

OpenStudy (anonymous):

let me think.........

OpenStudy (anonymous):

ok (1+50%)x=900 x=600 600*20%=120

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

how did you come up with that :)

OpenStudy (anonymous):

something wrong ?? ><

OpenStudy (anonymous):

i don't get the (1+50%)x=900 stuff

OpenStudy (anonymous):

i think it's totally wrong... it doesn't make sense -_-

OpenStudy (anonymous):

just know that profit is always relative to cost price.i.e the price at which you buy an item

OpenStudy (anonymous):

the price at which you sell it to some guy is your selling price

OpenStudy (anonymous):

why the price + % ? does it make sense? i didn't understand

OpenStudy (anonymous):

where is price + % ??

OpenStudy (anonymous):

cost price + percentage profit

OpenStudy (anonymous):

lets say cost price of a pup labrador =50$...now you want to sell it to a friend for 70$.If you do sell it..your profit will be 20$.your profit % will be 40%. to calculate profit %............. do 20=c.p*x/100. c.p=50%.so 20=50/100 *x..we get 20=x/2.=> x=40%

OpenStudy (anonymous):

@kryton1212

OpenStudy (anonymous):

ok